40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  TCS elog  Not logged in ELOG logo
Message ID: 110     Entry time: Thu Feb 24 10:23:31 2011
Author: Christopher Guido 
Type: Laser 
Category: Laser 
Subject: LTG initial noise 

Cheryl Vorvick, Chris Guido, Phil Willems

Attached is a PDF with some initial noise testing. There are 5 spectrum plots (not including the PreAmp spectrum) of the laser. The first two are with V_DC around 100 mV, and the other three are with V_DC around 200 mV. (As measured with the 100X gain preamplifier, so ideally 1 and 2 mV actual) We did one spectrum (at each power level)  with no attempt of noise reduction and one spectrum with the lights off and a make shift tent to reduce air flow. The 5th plot is at 200mv with the tent and the PZT on. (The other 4 have the PZT off).

 

The second plot is just the spectrums divided by their respectives V_DC to get an idea of the RIN.

Attachment 1: LTG_InitialTest.pdf  65 kB  | Hide | Hide all
LTG_InitialTest.pdf LTG_InitialTest.pdf
ELOG V3.1.3-